# A positive density of elliptic curves are diophantine stable in certain Galois Extensions

### v1

@@ 4.1. A suitable large family of elliptic curves
    (#¯$\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)=\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2}}\right)=\frac{(\ell^{2}-\ell)}{\ell^{2}}=1-\frac{1}{\ell}.$
@@
    #$\frac{\#\mathcal{F}_{\ell}}{\ell^{2n_{\ell}}}=\frac{1}{2^{21}}.$
@@
Let $\mathcal{F}_{\ell}^{\prime}$ be the complement of $\mathcal{F}_{\ell}$ in $\mathbb{Z}^{2}$. There is an absolute constant $C_{1}>0$ (independent of $\ell$) such that for all large enough values of $X$,
#$\#\mathcal{F}_{\ell}^{\prime}(X)\leq\frac{C_{1}}{\ell^{2}}X^{5/6}$
for all $\ell\notin T$.
@@
Assume without loss of generality that $\ell\geq 5$. We refer to the first part of the proof of Lemma 4.4. Thus, $\mathcal{F}_{\ell}^{\prime}$ is the set of $(A,B)\in\mathbb{Z}^{2}$ such that $\ell^{2}|(4A^{3}+27B^{2})$. We may write $\mathcal{F}_{\ell}^{\prime}=\mathcal{B}_{1}\cup\mathcal{B}_{2}\cup\mathcal{B}_{3}$, where
  • $\mathcal{B}_{1}$ consists of pairs $(A,B)$ for which $\ell|A$ and $\ell|B$,
  • $\mathcal{B}_{2}$ consists of pairs for which $\ell\nmid A,B$ and $-4A^{3}\equiv 27B^{2}\pmod{\ell^{2}}$, and $4A^{3}+27B^{2}\neq 0$.
  • Finally, let $\mathcal{B}_{3}$ consist of the pairs for which $04A^{3}+27B^{2}=$ 0.
@@
#$\begin{split}\#\mathcal{B}_{1}(X)=&\{(A,B)\in\mathcal{B}_{1}\mid\operatorname{max}\{|A|^{3},B^{2}\}\leq X\}\\ \leq&4\left\lfloor\frac{X^{1/3}}{\ell}\right\rfloor\left\lfloor\frac{X^{1/2}}{\ell}\right\rfloor\leq\frac{4X^{5/6}}{\ell^{2}}.\end{split}$
For each $A$ there are at most $2$ residue classes of $B$ modulo $\ell$ which satisfy $-4A^{3}\equiv 27B^{2}\pmod{\ell}$. Each residue class $B\pmod{\ell}$ lifts to a unique residue class modulo $\ell^{2}$ that satisfies $-4A^{3}\equiv 27B^{2}\pmod{\ell^{2}}$, by Hensel’s lemma. This requires that $B\not\equiv 0\pmod{\ell}$. The number of such $B$s in the interval $[-X^{1/2},X^{1/2}]$ in a given reside class modulo $\ell^{2}$ is $\frac{2X^{1/2}}{\ell^{2}}+O(1)$. Thus, in all, there are $2X^{1/3}(\frac{2X^{1/2}}{\ell^{2}}+O(1))$ values in #$\#\mathcal{B}_{2}(X)$. Hence, #$\#\mathcal{B}_{2}(X)\leq\frac{4X^{5/6}}{\ell^{2}}+O(X^{1/3})$. Finally, it is easy to see that #$\#\mathcal{B}_{3}(X)=o(X^{5/6})$ since the value of $A$ determines $B$ for $(A,B)\in\mathcal{B}_{3}(X)$. Thus, taking any constant $C_{1}>8$ will work. ∎
@@
#$\#\mathcal{F}_{\ell}(X)\sim 4\delta_{\ell}X^{5/6},$
where $\delta_{\ell}=\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2}}$. We estimate the size of $\#\overline{\mathcal{F}}_{\ell}$ for $\ell\in Z$.
@@
    #¯$\#\overline{\mathcal{F}}_{\ell}\geq\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-76}{\sqrt{\ell}}\right).$
@@
    #¯$\#\overline{\mathcal{F}}_{\ell}\geq\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-15.18}{\sqrt{\ell}}\right).$
@@
#$\#\mathcal{F}_{\ell}\geq\ell^{2}-\#\mathcal{A}_{1}-\#\mathcal{A}_{2}-\#\mathcal{A}_{3}.$
@@
#$\#\mathcal{A}_{2}=\#\mathcal{A}_{3}.$
As was discussed in the proof of Lemma 4.4, we have that #$\#\mathcal{A}_{1}=\ell$. Then it follows from [How93] that
#$\#\mathcal{A}_{2}\leq\left(\frac{\ell^{2}-\ell}{p-1}\right)\left(1+2.53\frac{p(p+1)}{\sqrt{\ell}}\right).$
@@
#$\#\mathcal{F}_{\ell}\geq\ell^{2}-\ell-2\left(\frac{\ell^{2}-\ell}{p-1}\right)\left(1+2.53\frac{p(p+1)}{\sqrt{\ell}}\right).$
@@
#$\#\mathcal{F}_{\ell}\geq\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-76}{\sqrt{\ell}}\right).$
@@
#$\begin{split}\#\mathcal{F}_{\ell}\geq&\ell^{2}-\#\mathcal{A}_{1}-\#\mathcal{A}_{2},\\ \geq&\ell^{2}-\ell-\left(\frac{\ell^{2}-\ell}{2}\right)\left(1+\frac{15.18}{\sqrt{\ell}}\right)\\ =&\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-15.18}{\sqrt{\ell}}\right).\end{split}$
@@
In fact, the code gives the exact values of $\#\overline{\mathcal{F}}_{\ell}$.
@@ 4.2. Proof of the main result
$\eta_{p}:=\begin{cases}\frac{1}{4}&\text{ if }p=3;\\ \frac{3}{8}&\text{ if }p=5;\end{cases}$
@@
Let $E$ be an elliptic curve in the set $\mathcal{E}^{0}$. By Proposition 3.14, the Assumption 2.1 is satisfied for $E$ (and $K=\mathbb{Q}$). It follows from Proposition 2.7 that $\operatorname{Sel}_{p}(E/L)=0$. In particular, the rank of $E(L)$ is 00. Let $S_{3}$ be the intersection $\mathcal{E}^{0}\cap S_{1}$. Given $E\in S_{3}$, we find that $\operatorname{rank}E(L)=0$ and $E(L)_{\operatorname{tors}}=0$, hence, $E(L)=0$.
@@ 4.3. An example
We illustrate Theorem 3.8 via an example. We fix $p=3$, and $L=\mathbb{Q}(\mu_{7})^{+}$ to be the maximal real subfield of $\mathbb{Q}(\mu_{7})$. Then $L/\mathbb{Q}$ is a cubic Galois extension with Galois group $G:=\operatorname{Gal}(L/K)\simeq\mathbb{Z}/3\mathbb{Z}$. The only prime that ramifies in $L$ is $7$, and thus $Z=\{7\}$ and $T=\{2,3,7\}$. Recall that $\mathfrak{A}_{7}$ is the number of $(\overline{A},\overline{B})\in(\mathbb{Z}/7\mathbb{Z})^{2}$ such that
@@
  • $E_{\overline{A},\overline{B}}(\mathbb{F}_{7})[3]=0$ and $E_{\overline{A},-\overline{B}}(\mathbb{F}_{7})[3]=0$.
Computations on Sagemath show that $\mathfrak{A}_{7}=12$. Thus, we find that the lower density of elliptic curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$ is
$\geq\eta_{3}\prod_{\ell}\delta_{\ell}=\frac{1}{4}\times\frac{1}{2^{21}}\times\left(1-\frac{1}{3}\right)\times\frac{12}{49}\times\prod_{\ell\neq 2,3,7}\left(1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}\right).$

### v2

@@ 4.1. A suitable large family of elliptic curves
    $\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)=\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2}}\right)=\frac{(\ell^{2}-\ell)}{\ell^{2}}=1-\frac{1}{\ell}.$
@@
    $\frac{\#\mathcal{F}_{\ell}}{\ell^{2n_{\ell}}}=\frac{1}{2^{21}}.$
@@
Let $z\geq 5$ be a real number $\mathcal{F}_{\ell}^{\prime}$ be the complement of $\mathcal{F}_{\ell}$ in $\mathbb{Z}^{2}$. There is an absolute constant $C_{1}>0$ (independent of $\ell$) such that for all large enough values of $X$,
$\left(\bigcup_{\ell>z}\mathcal{F}_{\ell}^{\prime}(X)\right)\leq\sum_{\ell>z}\frac{C_{1}}{\ell^{2}}X^{5/6}+o(X^{5/6}).$
@@
Assume without loss of generality that $\ell\geq 5$. We refer to the first part of the proof of Lemma 4.4. Thus, $\mathcal{G}_{z}=\bigcup_{\ell>z}\mathcal{F}_{\ell}^{\prime}$ is the set of $(A,B)\in\mathbb{Z}^{2}$ such that $\ell^{2}|(4A^{3}+27B^{2})$ for some prime $\ell>z$. We may write $\mathcal{G}_{z}=\mathcal{B}_{1}\cup\mathcal{B}_{2}\cup\mathcal{B}_{3}$, where
  • $\mathcal{B}_{1}$ consists of pairs $(A,B)$ for which $\ell|A$ and $\ell|B$ for some prime $\ell>z$,
  • $\mathcal{B}_{2}$ consists of pairs for which $\ell\nmid A,B$ and $4A^{3}\equiv 27B^{2}\pmod{\ell^{2}}$, and $4A^{3}+27B^{2}\neq 0$ for some prime $\ell>z$.
  • Finally, let $\mathcal{B}_{3}$ consist of the pairs for which $4A^{3}+27B^{2}=0$.
@@
$\begin{split}\#\mathcal{B}_{1}(X)=&\{(A,B)\in\mathcal{B}_{1}\mid\operatorname{max}\{|A|^{3},B^{2}\}\leq X\}\\ \leq&\sum_{\ell>z}4\left\lfloor\frac{X^{1/3}}{\ell}\right\rfloor\left\lfloor\frac{X^{1/2}}{\ell}\right\rfloor\leq\sum_{\ell>z}\frac{4X^{5/6}}{\ell^{2}}.\end{split}$
Now suppose that $(A,B)\in\mathcal{B}_{2}(X)$. Note that $|4A^{3}+27B^{2}|\leq 31X$ is nonzero and is divisible by $\ell^{2}$ for some $\ell>z$. Therefore, we find that $\ell\leq\sqrt{31X}$. For each $A$ there are at most $2$ residue classes of $B$ modulo $\ell$ which satisfy $4A^{3}\equiv 27B^{2}\pmod{\ell}$. Each residue class $B\pmod{\ell}$ lifts to a unique residue class modulo $\ell^{2}$ that satisfies $4A^{3}\equiv 27B^{2}\pmod{\ell^{2}}$, by Hensel’s lemma. This requires that $B\not\equiv 0\pmod{\ell}$. The number of such $B$s in the interval $X^{1/2},X^{1/2}]$ in a given residue class modulo $\ell^{2}$ is $\frac{2X^{1/2}}{\ell^{2}}+O(1)$. Thus, in all,
$\begin{split}\#\mathcal{B}_{2}(X)\leq&\sum_{z<\ell\leq\sqrt{31X}}2X^{1/3}\left(\frac{2X^{1/2}}{\ell^{2}}+O(1)\right),\\ \leq&\sum_{\ell>z}\frac{4X^{5/6}}{\ell^{2}}+O\left(X^{1/3}\pi(\sqrt{31X})\right),\\ =&\sum_{\ell>z}\frac{4X^{5/6}}{\ell^{2}}+O\left(\frac{X^{5/6}}{\log X}\right).\end{split}$
Hence, $\mathcal{B}_{2}(X)\leq\sum_{\ell>z}\frac{4X^{5/6}}{\ell^{2}}+o\left(X^{5/6}\right)$. Finally, it is easy to see that $\mathcal{B}_{3}(X)=o(X^{5/6})$ since the value of $A$ determines $B$ for $(A,B)\in\mathcal{B}_{3}(X)$. Thus, taking any constant $C_{1}\geq 8$ will work. ∎
@@
$\mathcal{F}_{\ell}(X)\sim 4\delta_{\ell}X^{5/6},$
where $\delta_{\ell}=\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2}}$. We estimate the size of $\overline{\mathcal{F}}_{\ell}$ for $\ell\in Z$.
@@
    $\overline{\mathcal{F}}_{\ell}\geq\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-76}{\sqrt{\ell}}\right).$
@@
    $\overline{\mathcal{F}}_{\ell}\geq\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-15.18}{\sqrt{\ell}}\right).$
@@
$\mathcal{F}_{\ell}\geq\ell^{2}-\#\mathcal{A}_{1}-\#\mathcal{A}_{2}-\#\mathcal{A}_{3}.$
@@
$\mathcal{A}_{2}=\#\mathcal{A}_{3}.$
As was discussed in the proof of Lemma 4.4, we have that $\mathcal{A}_{1}=\ell$. Then it follows from [Howe] that
$\mathcal{A}_{2}\leq\left(\frac{\ell^{2}-\ell}{p-1}\right)\left(1+2.53\frac{p(p+1)}{\sqrt{\ell}}\right).$
@@
$\mathcal{F}_{\ell}\geq\ell^{2}-\ell-2\left(\frac{\ell^{2}-\ell}{p-1}\right)\left(1+2.53\frac{p(p+1)}{\sqrt{\ell}}\right).$
@@
$\mathcal{F}_{\ell}\geq\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-76}{\sqrt{\ell}}\right).$
@@
$\begin{split}\#\mathcal{F}_{\ell}\geq&\ell^{2}-\#\mathcal{A}_{1}-\#\mathcal{A}_{2},\\ \geq&\ell^{2}-\ell-\left(\frac{\ell^{2}-\ell}{2}\right)\left(1+\frac{15.18}{\sqrt{\ell}}\right)\\ =&\left(\frac{\ell^{2}-\ell}{2}\right)\left(\frac{\sqrt{\ell}-15.18}{\sqrt{\ell}}\right).\end{split}$
@@
In fact, the code gives the exact values of $\overline{\mathcal{F}}_{\ell}$.
@@ 4.2. Proof of the main result
$\eta_{p}\mathrel{\mathop{\ordinarycolon}}=\begin{cases}\frac{1}{4}&\text{ if }p=3;\\ \frac{3}{8}&\text{ if }p=5;\end{cases}$
@@
Let $E$ be an elliptic curve in the set $\mathcal{E}^{0}$. By Proposition 3.14, the Assumption 2.1 is satisfied for $E$ (and $K=\mathbb{Q}$). It follows from Proposition 2.7 that $\operatorname{Sel}_{p}(E/L)=0$. In particular, the rank of $E(L)$ is 0. Let $S_{3}$ be the intersection $\mathcal{E}^{0}\cap S_{1}$. Given $E\in S_{3}$, we find that $\operatorname{rank}E(L)=0$ and $E(L)_{\operatorname{tors}}=0$, hence, $E(L)=0$.
@@ 4.3. An example
We illustrate Theorem 3.8 via an example. We fix $p=3$, and $L$ to be the cubic subfield of $\mathbb{Q}(\mu_{31})$. Note that $2$ splits in $L$. The only prime that ramifies in $L$ is $31$, and thus $Z=\{31\}$ and $T=\{2,3,31\}$. Recall that $\mathfrak{A}_{31}$ is the number of $(\overline{A},\overline{B})\in(\mathbb{Z}/31\mathbb{Z})^{2}$ such that
@@
  • $E_{\overline{A},\overline{B}}(\mathbb{F}_{31})[3]=0$ and $E_{\overline{A},-\overline{B}}(\mathbb{F}_{31})[3]=0$.
Computations on Sagemath below
flag = 0
for A in GF(31):
     for B in GF(31):
            if (4*A^3+27*B^2)%31 !=0:
                E=EllipticCurve([A,B])
                if (E.abelian_group().order())%3 != 0:
                    print("A=", A, " B=", B, " #E(A,B)=", E.order())
                    flag = flag +1
print(flag)
show that $\mathfrak{A}_{31}=585$. Thus, we find that the lower density of elliptic curves $E_{/\mathbb{Q}}$ that are diophantine stable in $L$ is
$\geq\eta_{3}\prod_{\ell}\delta_{\ell}=\frac{1}{4}\times\frac{1}{2^{21}}\times\left(1-\frac{1}{3}\right)\times\frac{585}{961}\times\prod_{\ell\neq 2,3,31}\left(1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}\right).$